# Periodic orbits of non-degenerate lacunary contact forms on prequantization bundles

### v2

@@ Abstract.
A non-degenerate contact form is lacunary if the indexes of every contractible periodic Reeb orbit have the same parity. To the best of our knowledge, every contact form with finitely many periodic orbits known so far is non-degenerate and lacunary. We show that every non-degenerate lacunary contact form on a suitable prequantization of a closed symplectic manifold $B$ has precisely $r_{B}$ contractible closed orbits, where $r_{B}=\dim{\mathrm{H}}_{*}(B;{\mathbb{Q}})$. Examples of such prequantizations include the standard contact sphere and the unit cosphere bundle of a compact rank one symmetric space (CROSS). We also consider some prequantizations of orbifolds, like lens spaces and the unit cosphere bundle of lens spaces, and obtain multiplicity results for these prequantizations.
@@
MA was partially funded by FCT/Portugal through UID/MAT/04459/2020. LM was partially supported by CNPq/Brazil.
@@ 1.1. Introduction and Main Results
A result in this direction is the following theorem proved by Ginzburg, Gürel and Macarini [27]. In order to state it, we need some preliminary definitions. Given a closed Reeb orbit $\gamma$, let ${\mu}(\gamma)$ and ${\hat{\mu}}(\gamma)$ be its Conley-Zehnder index and mean index respectively. A contact form $\alpha$ supporting $\xi$ is index-positive (resp. index-negative) if the mean index ${\hat{\mu}}(\gamma)$ is positive (resp. negative) for every contractible periodic orbit $\gamma$ of $\alpha$. We say that $\alpha$ is index-admissible if every contractible closed orbit $\gamma$ of $\alpha$ satisfies ${\mu}(\gamma)>3-n$. Given a symplectic manifold $B$, let
@@
The theorem in [27] is actually a bit more general and relaxes the assumption that ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$ if $c_{B}>n$. However, in this case the lower bound obtained in [27] is no longer given by the total rank of the homology of $B$. To the best of our knowledge, every prequantization of a symplectic manifold $B$ that admits a contact form with finitely many closed orbits known so far has the property that ${\mathrm{H}}_{*}(B;\mathbb{Q})$ is lacunary.
@@
Every CROSS $N$ admits a metric such that all of its geodesics are periodic with the same minimal period; in other words, the geodesic flow generates a free circle action on $S^{*}N$. (To the best of our knowledge, a CROSS is the only known example so far of a closed Riemannian manifold admitting such a metric [7].) Thus the unit cosphere bundle $S^{*}N$ is a prequantization of a closed symplectic manifold $(B,\omega)$. Moreover, a homological computation shows that ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$; see [43, page 141]. In this case, the total rank $r_{B}$ of ${\mathrm{H}}_{*}(B;\mathbb{Q})$ and the minimal Chern number $c_{B}$ are given by Table 1.
@@
$S^{*}S^{2}$ or $S^{*}{\mathbb{R}}P^{2}$ $2$ $2$
$S^{*}S^{m}$ or $S^{*}{\mathbb{R}}P^{m}$ with $m>2$ even $m$ $m-1$
$S^{*}S^{m}$ or $S^{*}{\mathbb{R}}P^{m}$ with $m$ odd $m+1$ $m-1$
@@
The answer is also positive in the case of $M=S^{3}$ with the standard contact structure as was recently proved by Cristofaro-Gardiner, Hryniewicz, Hutchings and Liu [16]. In higher dimensions, it is a widely open and very difficult problem.
[…]
@@
Let $(M^{2n+1},\xi)$ be a prequantization $S^{1}$-bundle of a closed symplectic manifold $(B,\omega)$ such that $\omega|_{\pi_{2}(B)}\neq 0$, $c_{B}>n/2$ and ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$. Assume that $B$ is spherically positive monotone, that is, $c_{1}(TB)|_{\pi_{2}(B)}=\lambda\omega|_{\pi_{2}(B)}$ for some $\lambda>0$. Let $\alpha$ be a non-degenerate lacunary contact form supporting $\xi$. Suppose that one of the following two conditions holds:
@@
    $c_{1}(\xi)|_{H_{2}(M,\mathbb{Q})}=0$ and $\alpha$ is index-admissible.
@@
The hypothesis (F) in the previous theorem dropped the assumption that the symplectic form on $W$ is aspherical in Theorem 1.1. This is possible due to the use of Novikov fields and an action filtration introduced by McLean and Ritter [41]. Furthermore, the hypothesis (NF) in the previous theorem is weaker than the assumption (NF) in Theorem 1.1: we allow $c_{1}(\xi)$ to be torsion. See Section 2.2. Hypotheses (F) and (NF) in Theorem 1.1 can be weakened as in the hypotheses of Theorem 1.6; see Sections 2 and 6.
@@
An example of a prequantization satisfying the conditions of Theorem 1.6 under the assumption (F) such that the symplectic form of the filling is not aspherical is given by the prequantization $M$ of $S^{2}\times S^{2}\times S^{2}$ endowed with the product symplectic form, where the symplectic form on $S^{2}$ has area 1. One can check that $M$ is simply connected and the filling $W$ of $M$, given by a $4$-ball bundle over $S^{2}\times S^{2}$, has vanishing first Chern class. However, the symplectic form is not aspherical since the zero section $S^{2}\times S^{2}$ is a symplectic submanifold.
@@
Let $(M^{2n+1},\xi)$ be a prequantization $S^{1}$-bundle of a closed symplectic manifold $(B,\omega)$ and $\alpha$ a non-degenerate lacunary contact form supporting $\xi$ as in the previous theorem. Assume that $\pi_{1}(M)$ is finite. Then $\alpha$ has precisely $r_{B}$ simple closed orbits, where $r_{B}=\dim{\mathrm{H}}_{*}(B;\mathbb{Q})$.
@@
The last corollary improves results due to Duan-Liu-Ren [18] and Duan-Xie [20]. In the case of the standard contact sphere $S^{2n+1}$, it follows easily from Theorem 1.1 and [32, Theorem 1.5]. (Actually, we do not need to use Theorem 1.1: we just have to note that a non-degenerate lacunary contact form on $S^{2n+1}$ is dynamically convex (i.e. every closed orbit has index at least $n+2$) and use the fact that a non-degenerate dynamically convex contact form on $S^{2n+1}$ has at least $n+1$ simple closed orbits [3, 33]. On the other hand, by [32, Theorem 1.5] the number of simple periodic orbits of a non-degenerate lacunary contact form on $S^{2n+1}$ is at most $n+1$.) As mentioned before, the unit disk bundle $D^{*}N$ meets the assumption (F) of Theorem 1.6 whenever the dimension of $N$ is bigger than two. If $N$ is a surface then it is given by either $S^{2}$ or ${\mathbb{R}}P^{2}$ and in this case the result follows easily from Theorem 1.1 and [32, Theorem 1.5].
@@
Let $(M^{2n+1},\xi)$ be a prequantization $S^{1}$-bundle of a closed symplectic manifold $(B,\omega)$ such that $\omega|_{\pi_{2}(B)}\neq 0$, $c_{B}>n/2$ and ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$. Assume that $B$ is spherically positive monotone, that is, $c_{1}(TB)|_{\pi_{2}(B)}=\lambda\omega|_{\pi_{2}(B)}$ for some $\lambda>0$ and that $M$ satisfies the hypothesis (F) of Theorem 1.6. Suppose that $B$ admits a Hamiltonian $T^{d}$-action with isolated fixed points and consider the corresponding lifted $T^{d+1}$-action on $M$. Let $G$ be a finite subgroup of $T^{d+1}$ acting freely on $M$ and let ${\widebar{M}}=M/G$ be the quotient. Then every non-degenerate lacunary contact form $\alpha$ on ${\widebar{M}}$ has precisely $r_{B}$ simple contractible closed orbits, where $r_{B}=\dim{\mathrm{H}}_{*}(B;\mathbb{Q})$. Moreover, the lifts of these orbits to $M$ are symmetric closed orbits of the lifted contact form ${\hat{\alpha}}$ on $M$ with respect to the $G$-action.
@@
The main point in the previous theorem is that ${\widebar{M}}$ is not, in general, a prequantization of a symplectic manifold: it is the prequantization of a symplectic orbifold. (Indeed, the $S^{1}$-action on ${\widebar{M}}$ is not free in general: it is only locally free.) For instance, symplectic manifolds meeting the assumptions of the previous theorem are monotone toric closed symplectic manifolds $B$ such that $c_{B}>n/2$. Take the particular case where $B={\mathbb{C}}P^{n}$ with the standard symplectic form $\omega$ normalized so that $\omega$ evaluated at a generator of ${\mathrm{H}}_{2}(B,\mathbb{Z})\cong\mathbb{Z}$ is ±$\pm 1$. Then $M$ is the standard contact sphere $S^{2n+1}$ which of course has a filling $W$ satisfying the hypotheses of the theorem. Clearly, ${\mathbb{C}}P^{n}$ has a Hamiltonian $T^{n}$-action with isolated fixed points. The lifted $T^{n+1}$-action on $S^{2n+1}$, regarded as a subset of $\mathbb{C}^{n+1}$, is given by
$(\theta_{0},\dots,\theta_{n})\cdot(z_{0},\dots,z_{n})\mapsto(e^{i\theta_{0}}z_{0},\dots,e^{i\theta_{n}}z_{n}).$ (1.1)
@@
$\psi(z_{0},\dots,z_{n})=\left(e^{\frac{2\pi i\ell_{0}}{p}}z_{0},e^{\frac{2\pi i\ell_{1}}{p}}z_{1},\dots,e^{\frac{2\pi i\ell_{n}}{p}}z_{n}\right),$ (1.2)
@@
This group has a maximal torus $T$ of dimension $\lfloor\frac{m+1}{2}\rfloor$. Suppose now that $m=2d-1$ is odd so that the dimension of $T$ is $d$. The corresponding action of $T^{d}$ on $S^{m}\subset\mathbb{C}^{d}$ is given by rotations in each coordinate as in (1.1). The lifted action of $T^{d}$ to $S^{*}S^{m}$ coincides with the lift of the Hamiltonian $T^{d}$-action on $B$ to $S^{*}S^{m}$ which commutes with the $S^{1}$-action induced by the geodesic flow, generating the aforementioned contact action of $T^{d+1}$ on $S^{*}S^{m}$. Given an integer $p\geq 1$, consider the $\mathbb{Z}_{p}$-action on $S^{m}$ generated by the map $\psi$ as in (1.2) with the weights $\ell_{0},\ldots,\ell_{d-1}$ coprime with $p$. The quotient $S^{m}/\mathbb{Z}_{p}$ is the lens space $L^{m}_{p}(\ell_{0},\dots,\ell_{d-1})$ and the quotient $S^{*}S^{m}/\mathbb{Z}_{p}$ with respect to the lifted action is the unit cosphere bundle of this lens space. Therefore, we conclude the following corollary.
@@
It follows from the previous discussion and the proof of Theorem 1.13, presented in Section 5, that a lens space admits a Finsler metric with finitely many simple closed orbits (more precisely, $m+1$ closed orbits if the lens space has dimension $m$); see Remark 5.4. The point here is that the $\mathbb{Z}_{p}$-action on $S^{*}S^{m}$ is the lift of an action on $S^{m}$ which is isometric with respect to the Katok-Ziller metric. Thus, a CROSS is not the only example of a closed manifold admitting a Finsler metric with finitely many simple closed orbits; cf. [43, Page 140].
@@
Theorem 1.13 raises the question if Theorem 1.6 can be generalized to prequantizations of orbifolds. Morever, another natural question is if we can relax the assumption on the minimal Chern number $c_{B}$. We plan to work on it in a forthcoming paper.

### v3

@@ Abstract.
A non-degenerate contact form is lacunary if the indexes of every contractible periodic Reeb orbit have the same parity. To the best of our knowledge, every contact form with finitely many periodic orbits known so far is non-degenerate and lacunary. We show that every non-degenerate lacunary contact form on a suitable prequantization of a closed symplectic manifold $B$ has precisely $r_{B}$ contractible closed orbits, where $r_{B}=\dim{\mathrm{H}}_{*}(B;{\mathbb{Q}})$. Examples of such prequantizations include the standard contact sphere, the unit cosphere bundle of a compact rank one symmetric space (CROSS) and many others. We also consider some prequantizations of orbifolds, like lens spaces and the unit cosphere bundle of lens spaces, and obtain multiplicity results for these prequantizations.
@@
MA was partially funded by FCT/Portugal through UID/MAT/04459/2020. LM was partially supported by FAPERJ and CNPq/Brazil.
@@ 1.1. Introduction and Main Results
A result in this direction is the following theorem proved by Ginzburg, Gürel and Macarini [30]. In order to state it, we need some preliminary definitions. A closed Reeb orbit $\gamma$ is non-degenerate if its linearized Poincaré map does not have eigenvalue one. A contact form $\alpha$ is non-degenerate if every closed orbit $\gamma$ of $\alpha$ is non-degenerate. Given a closed Reeb orbit $\gamma$, let ${\mu}(\gamma)$ and ${\hat{\mu}}(\gamma)$ be its Conley-Zehnder index and mean index respectively (see, for instance, [45]). A contact form $\alpha$ supporting $\xi$ is index-positive (resp. index-negative) if the mean index ${\hat{\mu}}(\gamma)$ is positive (resp. negative) for every contractible periodic orbit $\gamma$ of $\alpha$. We say that $\alpha$ is index-admissible if every contractible closed orbit $\gamma$ of $\alpha$ satisfies ${\mu}(\gamma)>3-n$. Given a symplectic manifold $B$, let
@@
The theorem in [30] (namely, [30, Theorem 2.1]) is actually a bit more general and relaxes the assumption that ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$ if $c_{B}>n$. However, in this case the lower bound obtained in [30] is no longer given by the total rank of the homology of $B$. To the best of our knowledge, every prequantization of a symplectic manifold $B$ that admits a contact form with finitely many closed orbits known so far has the property that ${\mathrm{H}}_{*}(B;\mathbb{Q})$ is lacunary.
@@
Every CROSS $N$ admits a metric such that all of its geodesics are periodic with the same minimal period; in other words, the geodesic flow generates a free circle action on $S^{*}N$. (To the best of our knowledge, a CROSS is the only known example so far of a closed manifold admitting such a metric [8].) Thus the unit cosphere bundle $S^{*}N$ is a prequantization of a closed symplectic manifold $(B,\omega)$. Moreover, a homological computation shows that ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$; see [47, page 141]. In this case, the total rank $r_{B}$ of ${\mathrm{H}}_{*}(B;\mathbb{Q})$ and the minimal Chern number $c_{B}$ are given by Table 1.
@@
$S^{*}S^{2}$ or $S^{*}{\mathbb{R}}P^{2}$ 22 22
$S^{*}S^{m}$ or $S^{*}{\mathbb{R}}P^{m}$ with $m>2$ even $m$ m1m-1
$S^{*}S^{m}$ or $S^{*}{\mathbb{R}}P^{m}$ with $m$ odd $m+1$ m1m-1
@@
The answer is also positive in the case of $M=S^{3}$ with the standard contact structure as was recently proved by Cristofaro-Gardiner, Hryniewicz, Hutchings and Liu [18]. In higher dimensions, it is a widely open and very difficult problem; see however the recent work of Cineli-Ginzburg-Gürel [14] with important results on $S^{2n+1}$ under assumptions of symmetry, convexity and non-degeneracy of the contact form.
A periodic orbit is elliptic if every eigenvalue of its linearized Poincaré map has modulus one. All the examples of contact forms with finitely many closed orbits that we know so far are non-degenerate and have the property that every periodic orbit is elliptic. It is well known that if a non-degenerate periodic orbit $\gamma$ is elliptic then ${\mu}(\gamma)=n\ \text{(mod 2)}$. In particular, these contacts forms are geometrically perfect, that is, the indexes of every periodic orbit in the same free homotopy class have the same parity. Another example of geometrically perfect contact forms are those with Anosov Reeb flows (e.g. the geodesic flow of a metric with negative sectional curvature) [41]. In these examples, every periodic orbit is hyperbolic. We do not know examples of geometrically perfect contact forms with elliptic and hyperbolic orbits.
@@
Let $(M^{2n+1},\xi)$ be a prequantization $S^{1}$-bundle of a closed symplectic manifold $(B,\omega)$ such that $\omega|_{\pi_{2}(B)}\neq 0$ and ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$. Assume that $B$ is spherically positive monotone, that is, $c_{1}(TB)|_{\pi_{2}(B)}=\lambda\omega|_{\pi_{2}(B)}$ for some $\lambda>0$. Let $\alpha$ be a non-degenerate lacunary contact form supporting $\xi$. Suppose that one of the following two conditions holds:
@@
    $c_{1}(\xi)|_{H_{2}(M,\mathbb{Q})}=0$, $c_{B}>1$ and $\alpha$ is index-admissible.
@@
We originally proved this theorem under the assumption that $c_{B}>n/2$ as in Theorem 1.1. Indeed, the proof given in Section 4 assumes this hypothesis. In Appendix A, written jointly with Baptiste Serraille, we dropped completely this assumption.
@@
Instead of considering only contractible orbits, we can consider every periodic orbit. Then the assumption that the map $\pi_{1}(M)\to\pi_{1}(W)$ is injective in hypothesis (F) can be dropped. The price, of course, is that in this case the condition to be lacunary is stronger, since we ask that the indexes of every closed orbit have the same parity instead of just the indexes of the contractible ones. We consider this situation in [6].
@@
The hypothesis (F) in the previous theorem dropped the assumption that the symplectic form on $W$ is aspherical in Theorem 1.1. This is possible due to the use of Novikov fields and an action filtration introduced by McLean and Ritter [44]. Furthermore, the hypothesis (NF) in the previous theorem is weaker than the assumption (NF) in Theorem 1.1: we allow $c_{1}(\xi)$ to be torsion. See Sections 2 and 6.
Remark 1.9.
The assumption $c_{B}>1$ in (NF) is used to compute ${\mathrm{HC}}^{0}_{*}(M)$ when we do not have a filling in Proposition 2.1. More precisely, it is used to obtain an index-admissible contact form given by a small perturbation of the connection form; cf. Remark 1.18.
Remark 1.10.
An example of a prequantization satisfying the conditions of Theorem 1.6 under the assumption (F) such that the symplectic form of the filling is not aspherical is given by the prequantization $M$ of $S^{2}\times S^{2}\times S^{2}$ endowed with the product symplectic form, where the symplectic form on $S^{2}$ has area 1. One can check that $M$ is simply connected and has a filling $W$, given by a $4$-ball bundle over $S^{2}\times S^{2}$, with vanishing first Chern class. However, the symplectic form on $W$ is not aspherical since the zero section $S^{2}\times S^{2}$ is a symplectic submanifold. For more examples, see Remark 1.15.
@@
Let $(M^{2n+1},\xi)$ be a prequantization $S^{1}$-bundle of a closed symplectic manifold $(B,\omega)$ and $\alpha$ a non-degenerate lacunary contact form supporting $\xi$ as in Theorem 1.6. Assume that $\pi_{1}(M)$ is finite. Then $\alpha$ has precisely $r_{B}$ simple closed orbits, where $r_{B}=\dim{\mathrm{H}}_{*}(B;\mathbb{Q})$.
@@
The last corollary improves results due to Duan-Liu-Ren [21] and Duan-Xie [23]. In the case of the standard contact sphere $S^{2n+1}$, it follows easily from Theorem 1.1 and [35, Theorem 1.5]. (Actually, we do not need to use Theorem 1.1: we just have to note that a non-degenerate lacunary contact form on $S^{2n+1}$ is dynamically convex (i.e. every closed orbit has index at least $n+2$) and use the fact that a non-degenerate dynamically convex contact form on $S^{2n+1}$ has at least $n+1$ simple closed orbits [3, 36]. On the other hand, by [35, Theorem 1.5] the number of simple periodic orbits of a non-degenerate lacunary contact form on $S^{2n+1}$ is at most $n+1$.) As mentioned before, the unit disk bundle $D^{*}N$ meets the assumption (F) of Theorem 1.6 whenever the dimension of $N$ is bigger than two. If $N$ is a surface then it is given by either $S^{2}$ or ${\mathbb{R}}P^{2}$ and in this case the result follows easily from Theorem 1.1 and [35, Theorem 1.5].
Remark 1.15 (More examples).
As explained before, we proved originally Theorem 1.6 under the assumption that $c_{B}>n/2$ which is satisfied by the examples of the previous corollary. Without this hypothesis, we have plenty of other examples of prequantizations satisfying the assumptions of Theorem 1.6. As a matter of fact, let us consider first examples meeting the assumption (F). We have, for instance, that every closed toric symplectic manifold $(B,\omega)$ satisfies $\omega|_{\pi_{2}(B)}\neq 0$ and ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$. We claim that if $B$ is monotone and has dimension four then it admits a number $m\in\mathbb{N}$ such that $\omega/m]\in{\mathrm{H}}^{2}(B,\mathbb{Z})$ and the prequantization $(M,\xi)$ of $(B,\omega/m)$ has a filling as in hypothesis (F). Indeed, take $m\in\mathbb{N}$ such that $\omega/m]\in{\mathrm{H}}^{2}(B,\mathbb{Z})$ and $\omega/m]$ is a primitive class, that is, there is no $c\in{\mathrm{H}}^{2}(B,\mathbb{Z})$ such that $\omega/m]=lc$ for some integer $l\geq 2$. This last condition is equivalent to the existence of an element $S\in\pi_{2}(B)\cong H_{2}(B,\mathbb{Z})$ such that $\langle[\omega/m],S\rangle=1$. Then one can easily check, using the long exact homotopy sequence of the $S^{1}$-bundle $M\to B$, that $M$ is simply connected (note that $B$ is simply connected because it is toric). Moreover, since $B$ is monotone, we have that $c_{1}(\xi)$ is torsion; cf. Remark 1.11. But, since $M$ is simply connected, we conclude from the universal coefficient theorem for cohomology that ${\mathrm{H}}^{2}(M,\mathbb{Z})$ is torsion-free and hence $c_{1}(\xi)=0$. Thus, $(M,\xi)$ is a simply connected toric contact manifold with $c_{1}(\xi)=0$. It can be proved that when $M$ has dimension five then it admits a (toric) symplectic filling $W$ such that $c_{1}(TW)=0$; see [4, Section 5]. (Notice that the symplectic form on $W$ is not aspherical in general.) Hence, $W$ satisfies hypothesis (F). Finally, note that if $B$ is monotone then it is positive monotone because it is toric. We can easily construct several examples in higher dimensions; cf. [4, Section 5]. Examples satisfying (NF) are given by any prequantization of a closed monotone toric symplectic manifold with minimal Chern number bigger than one (e.g. monotone products of complex projective spaces).
@@
Let $(M^{2n+1},\xi)$ be a prequantization $S^{1}$-bundle of a closed symplectic manifold $(B,\omega)$ such that $\omega|_{\pi_{2}(B)}\neq 0$, $c_{B}>1$ and ${\mathrm{H}}_{k}(B;\mathbb{Q})=0$ for every odd $k$. Assume that $B$ is spherically positive monotone, that is, $c_{1}(TB)|_{\pi_{2}(B)}=\lambda\omega|_{\pi_{2}(B)}$ for some $\lambda>0$ and that $M$ satisfies the hypothesis (F) of Theorem 1.6. Suppose that $B$ admits a Hamiltonian $T^{d}$-action with isolated fixed points and consider the corresponding lifted $T^{d+1}$-action on $M$. Let $G$ be a finite subgroup of $T^{d+1}$ acting freely on $M$ and let ${\widebar{M}}=M/G$ be the quotient. Then every non-degenerate lacunary contact form $\alpha$ on $\widebar{M}$} has precisely $r_{B}$ simple contractible closed orbits, where $r_{B}=\dim{\mathrm{H}}_{*}(B;\mathbb{Q})$. Moreover, the lifts of these orbits to $M$ are symmetric closed orbits of the lifted contact form $^{\hat{\alpha}}$ on $M$ with respect to the $G$-action.
@@
Remark 1.18.
The hypothesis that $c_{B}>1$ is used only to ensure that $\alpha$ is index-admissible and therefore we can argue as in Theorem 1.6 under the assumption (NF); cf. Remark 1.9. This hypothesis can be dropped once $\widebar{M}$} has a filling satisfying (F).
The main point in the previous theorem is that $\widebar{M}$} is not, in general, a prequantization of a symplectic manifold: it is the prequantization of a symplectic orbifold. (Indeed, the $S^{1}$-action on $\widebar{M}$} is not free in general: it is only locally free.) For instance, symplectic manifolds meeting the assumptions of the previous theorem are monotone toric closed symplectic manifolds $B$ such that $c_{B}>1$. Take the particular case where $B={\mathbb{C}}P^{n}$ with the standard symplectic form $\omega$ normalized so that $\omega$ evaluated at a generator of ${\mathrm{H}}_{2}(B,\mathbb{Z})\cong\mathbb{Z}$ is $\pm 1$. Then $M$ is the standard contact sphere $S^{2n+1}$ which of course has a filling $W$ satisfying the hypotheses of the theorem. Clearly, ${\mathbb{C}}P^{n}$ has a Hamiltonian $T^{n}$-action with isolated fixed points. The lifted $T^{n+1}$-action on $S^{2n+1}$, regarded as a subset of $\mathbb{C}^{n+1}$, is given by
(1.1) $(\theta_{0},\dots,\theta_{n})\cdot(z_{0},\dots,z_{n})\mapsto(e^{i\theta_{0}}z_{0},\dots,e^{i\theta_{n}}z_{n}).$
@@
(1.2) $\psi(z_{0},\dots,z_{n})=\left(e^{\frac{2\pi i\ell_{0}}{p}}z_{0},e^{\frac{2\pi i\ell_{1}}{p}}z_{1},\dots,e^{\frac{2\pi i\ell_{n}}{p}}z_{n}\right),$
@@
This group has a maximal torus $T$ of dimension $\lfloor\frac{m+1}{2}\rfloor$. Suppose now that m=2d1m=2d-1 is odd so that the dimension of $T$ is $d$. The corresponding action of $T^{d}$ on $S^{m}\subset\mathbb{C}^{d}$ is given by rotations in each coordinate as in (1.1). The lifted action of $T^{d}$ to $S^{*}S^{m}$ coincides with the lift of the Hamiltonian $T^{d}$-action on $B$ to $S^{*}S^{m}$ which commutes with the $S^{1}$-action induced by the geodesic flow, generating the aforementioned contact action of $T^{d+1}$ on $S^{*}S^{m}$. Given an integer $p\geq 1$, consider the $\mathbb{Z}_{p}$-action on $S^{m}$ generated by the map $\psi$ as in (1.2) with the weights $\ell_{0},\ldots,\ell_{d-1}$ coprime with $p$. The quotient $S^{m}/\mathbb{Z}_{p}$ is the lens space $L^{m}_{p}(\ell_{0},\dots,\ell_{d-1})$ and the quotient $S^{*}S^{m}/\mathbb{Z}_{p}$ with respect to the lifted action is the unit cosphere bundle of this lens space. Therefore, we conclude the following corollary.
@@
It follows from the previous discussion and the proof of Theorem 1.16, presented in Section 5, that a lens space admits a Finsler metric with finitely many simple closed orbits (more precisely, $m+1$ closed orbits if the lens space has dimension $m$); see Remark 5.4. The point here is that the $\mathbb{Z}_{p}$-action on $S^{*}S^{m}$ is the lift of an action on $S^{m}$ which is isometric with respect to the Katok-Ziller metric. Thus, a CROSS is not the only example of a closed manifold admitting a Finsler metric with finitely many simple closed orbits; cf. [47, Page 140].
@@
Remark 1.23.
Similarly to Theorem 1.6, we originally proved Theorem 1.16 under the assumption that $c_{B}>n/2$. As in Remark 1.15, under the weaker hypothesis that $c_{B}>1$ we can obtain several new examples where we can apply Theorem 1.16.
Theorem 1.16 raises the question if Theorem 1.6 can be generalized to prequantizations of orbifolds. In [6], we obtain this generalization for several such prequantizations, including a large class of toric contact manifolds. We also prove multiplicity results like Theorem 1.1 for these prequantizations. In particular, we drop the assumption on the minimal Chern number in Theorem 1.1.